If (x+y)=3(x-y)
then find the value of 3xy÷2(x2-y2)

Grade:Select Grade

3 Answers

Ravi
9 years ago
Let x-y=u
x+y=v

So, u=3v...(I)
Also, (x2-y2)=ab
xy=(v2-u2)/4...(II)

Use i and II in the given expression to get the value of the expression.
.....................................................................................................................................
Ravi
9 years ago
Let x-y=u
x+y=v

So, u=3v...(I)
Also, (x2-y2)=ab
xy=(v2-u2)/4...(II)

Use i and II in the given expression to get the value of the expression.
.....................................................................................................................................
Ravi
9 years ago
Let x-y=u
x+y=v

So, u=3v...(I)
Also, (x2-y2)=ab
xy=(v2-u2)/4...(II)

Use i and II in the given expression to get the value of the expression.
​.....................................................................................................................................